### v1

@@ 4.4 DVFL for Split Training of Neural Networks
9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h,i},\nabla_{\theta_{h,i}}\mathcal{L}_{h,i})${Update host model}
@@
5 Experiments
As a toy example, we consider a system of $|\mathcal{G}|$ guests which can each see $\frac{1}{|\mathcal{G}|}^{th}$ of an handwritten digit from the MNIST dataset. For our experiments, we set $|\mathcal{G}|=4$ i.e. $g_{1}$’s dataset contains the top $28\times 7$ pixel patch of an MNIST image, $g_{2}$’s dataset contains the second , $g_{3}$’s contains the third, and $g_{4}$’s contains the bottom $28\times 7$ patch (see Fig. 3).
Figure 3: For our experiments, the input received by each guest for a given sample, as well as the reconstruction its autoencoder model produces.
We use a learnable decoder head and MSE reconstruction as an unsupervised objective.
Baseline
DVFL is an alternate strategy to VFL. Naturally we compare DVFL for ANNs to SplitNN, a direct application of VFL to ANNs. Naive VFL implicitly employs the “wait” protocol discussed in Sec 2.4.1. There are some simple additions to SplitNN that would allow a more meaningful comparison. Firstly, if the host sets a deadline, we may continue training despite missing inputs by skipping (we call this SplitNN-skip) or by supplying replacement values. For replacement values, we try a zero tensor (SplitNN-zeros). Hosts may ask guests to send the identities of the entity to which each of their messages belong. This way, a late message cannot be believed to be intended for the next iteration. Results on more baselines and datasets are presented in Appendix A. The choice of baseline is discussed in Appendix C.
5.1 Fault Simulation
Six hyperparameters control the fault model described in Section 2.4.1. $R^{(d)}_{i}$ is the fault rate for a live $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host}\}$. $R^{(u)}_{i}$ is the chance a dead $i$ rejoins the network before the next request. We discuss how exactly the faults are simulated in Appendix B.2. We assume no faults during label supervision or inference.
5.2 Experimental Setup
Table 1: We compare DVFL to standard VFL. Here, we report the median and twice the standard deviation of MNIST test performance when trained with various entities $i\in\{\mathrm{connection,guest,host}\}$ susceptible to fail at a rate of $R^{(d)}_{i}=0.3$. A failed resource has $R^{(u)}_{i}=\mathbb{P}_{\mathrm{rejoin}}\in\{1,0.5,0.1\}$ to be available next time it is requested. SplitNN, the baseline VFL algorithm, cannot train at all when there are any faults. On the other hand, our DVFL implementation degrades gracefully with faults. Using an explicit fault handling strategy within VFL does allow it to train. But the degradation with faults is sharper than DVFL and the variance in model performance is higher. Under fault-free circumstances, SplitNN slightly outperforms our DVFL implementation. With minor adjustments, SplitNN can tolerate some faults, but our approach is more resilient and stable.
Strategy Test Accuracy (%) with 0.3 Loss Rate
No Faults Communication Loss Guest Loss Host Loss
DVFL-NN 97.30$\pm$0.08 97.24$\pm$0.10 97.19$\pm$0.14 97.03$\pm$0.17 97.11$\pm$0.11 97.08$\pm$0.16 96.87$\pm$0.20 97.27$\pm$0.12 97.24$\pm$0.14 97.02$\pm$0.33
VFL (SplitNN) 97.80$\pm$0.17 N/A N/A N/A N/A N/A N/A N/A N/A N/A
VFL (SplitNN-skip) 97.84$\pm$0.16 97.50$\pm$0.54 96.60$\pm$0.51 77.21$\pm$5.34 97.44$\pm$0.57 96.65$\pm$0.45 80.22$\pm$4.31 97.65$\pm$0.41 97.59$\pm$0.50 97.00$\pm$0.89
VFL (SplitNN-zeros) 97.79$\pm$0.15 97.64$\pm$0.38 97.53$\pm$0.40 96.89$\pm$0.36 97.68$\pm$0.40 97.59$\pm$0.40 96.45$\pm$0.50 97.68$\pm$0.30 97.65$\pm$0.37 97.01$\pm$0.93
Performance under Faults
In order to show that it is possible to use DVFL even with crash faults, we train DVFL with $|\mathcal{H}|=4$ hosts under $R^{(d)}_{i}=0.3,R^{(d)}_{j}=0$ where $i\in\{\mathrm{connection},\mathrm{guest},\mathrm{host},\phi\},j\neq i$. Given a $R^{(d)}_{i}\neq 0$, we try $R^{(d)}_{i}\in\{0.1,0.5,1.0\}$.
Effect of Redundancy
In order to show that redundancy in hosts improves model performance and fault tolerance, we simulate DVFL training with $|\mathcal{H}|$ = 1, 2, 3, and 4 hosts. We measure test accuracy given $R_{\mathrm{connection}}\in\{0,0.3,0.6\}$ during training.
Learning from a Limited Intersection
We first fix a small intersection size, where all members of the intersection are labeled. We generate the remaining dataset for each device by shuffling the out-of-intersection samples and taking windows such that the remaining sample space is distributed evenly across all four guests. For example, if we wish to use $1024$ labeled samples, we shuffle samples indexed $1024-59999$ and split the $58976$ samples evenly, such that each guest gets $14744$ unique MNIST patches + $1024$ intersecting MNIST patches = $15768$ total samples. We try this experiment with $128$, $256$, $512$ and $1024$ labeled samples. We compare with SplitNN trained on only the labeled samples.
6 Results and Discussion
[…]
Figure 4: Model performance improves with redundancy, especially when there is a nonzero crash rate.
6.1 Graceful Degradation with Faults
We observe (Table 1) that model performance degrades with faults, but not catastrophically (unlike VFL with SplitNN). In ideal conditions, DVFL performance is slightly poorer than SplitNN. This is not unexpected, since our models are shallow. Wang et al. (2021) show performance improves with the length of backpropagation in blockwise-trained networks, and suggest it is due to increasing mutual information between shallower and deeper layers. Further, given a model, the reconstruction loss guides the feature extractor to features that will be useful for reconstruction– not necessarily good for classification. It is expected that the gap between DVFL and SplitNN will reduce with more general loss functions van den Oord et al. (2019); Chen et al. (2020b); Caron et al. (2020); Zbontar et al. (2021) and with self-supervised classification objectives. Further, observe that the degradation with faults in DVFL is much more graceful than our SplitNN variants. Not only does the SplitNN-wait/skip/stale performance drop more sharply, the variance increases more than it does with DVFL.
Guest faults are the worst
In the case of DVFL, the most severe type of fault is the guest loss. For example, when $R^{(d)}_{\mathrm{guest}}=0.3,R^{(u)}_{\mathrm{guest}}=0.1$ the accuracy drops from 97.30% to 96.87% while with communication and host loss it is 97.03 and 97.02% respectively. The performance of the combined model depends on unique outputs from guests– there is no redundancy. With a host or communication fault, guests are still able to update their own models (and there might be redundant hosts, too). In the case of SplitNN-skip and SplitNN-host, a communication fault is just as bad– a guest cannot send forward passes to (or receive backward passes from) the single host whether it or the connection is down. A host fault is equivalent to skipping that round for training.
Tolerating some faults pays off
Induced by the artificial input, the Dropout effect Srivastava et al. (2014) leads to a reduction in bias. This competes against the loss of information in the model. In the case of DVFL, models do not accept feedback, so the scope for information loss is smaller. In the case of SplitNN with missing input imputation, models do not respond to faults as well as DVFL. Still, the test accuracy degrades somwewhat gracefully for this reason.
Other failure modes
In this study, we only examine loss faults. Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults. For the forward pass, the first two might be mitigated within VFL itself. Since hosts request encoded features from a particular sample or batch of samples, they may ask guests to send back the identity of those samples along with the features (as we assumed in the variants of SplitNN. While the privacy is dubious, there is precedent in VFL: see Sun et al. (2021); Yang et al. (2023); Jiang et al. (2022). But, in order to correctly serialize, the host must maintain a cache of unused messages. For the backwad pass, if a guest has sent two outputs before receiving a single gradient, it has no way of knowing which output it corresponds to. DVFL is immune to these issues, since guests and hosts train independently to each other.
6.2 Redundancy Improves Performance
We observe (Figure 4) that test accuracy improves with the number of hosts with all other system parameters fixed. This is not surprising, as the owner’s aggregation of host results can be seen as a form of bagging Breiman (1996). As a corollary, the degradation of performance with communication faults is more graceful with increasing redundancy.
The role of the DVFL host
Observe that host faults do not hurt DVFL performance very much. DVFL hosts’ primary function is aggregation. A DVFL host can achieve this without any learnable parameters, by simply concatenating inputs and passing them onto the owner when requested. The other role a DVFL host might play is to project the concatenated host output to a smaller dimension. What the host model must learn is to preserve useful features in the new dimension.
Strategy Test Accuracy (%)
# of Labeled Samples
128 256 512 1024
DVFL 76.45 82.51 85.18 88.90
VFL (SplitNN) 65.07 77.51 83.09 87.26
6.3 Out-of-Intersection Training
Although VFL with SplitNN outperforms our DVFL ANN algorithm in our experiments with perfect conditions, DVFL significantly outperforms VFL on a limited labeled intersection. DVFL’s advantage is greater with fewer labels (Table 2).
6.4 Security & Privacy
DVFL achieves these results while also completely eliminating the possibility of gradient-based inference attacks. As reviewed in 3, complex mechanisms can be developed on top of SplitNN to provide privacy guarantees within some bounds. DVFL achieves complete gradient privacy without any overhead or model convergence issues.
7 Conclusion and Future Work
We present Decoupled Vertical Federated Learning, a strategy for joint learning on vertically partitioned data. Instead of adding complexity to cope with the data leakage and synchronization associated with BP, we eschew feedback altogether. Combined with the ability to reuse feedforward signals, this allows for fault tolerance and privacy in training. Training may continue even when there are faults, or there are a limited number of samples common to all guests.
DVFL also provides a high degree of privacy. Malicious or compromised guests can no longer use gradient feedback to figure out possibly sensitive information. Existing techniques to enforce privacy guarantees either hamper model convergence, are expensive to implement or both.
Our results suggest its performance to be comparable to SplitNN on various machine learning tasks. Unlike SplitNN, DVFL degrades gracefully with various kinds of loss faults. Even with explicit fault handling strategies, SplitNN’s performance drops off more sharply than DVFL’s. In general, DVFL’s asynchronicity provides a great deal of engineering flexibility. Some of it is discussed in the Appendix.
The next step for DVFL is to generalize it for more tasks and general architectures. VFL is readily compatible with a variety of machine learning models (including ensemble models). With a judicious choice of training objective for guest models, DVFL may also be extended past ANNs.
Another possible direction for DVFL is the introduction of more elaborate fault handling strategies. In this work, the aggregating “host” uses the the latest activation received from a “guest” client. This allows for a simple implementation using single-reader single-writer registers. We observe that this leads to meaningful representation learning– empirically, we only observe a 0.26% drop in accuracy on MNIST even when there is a 30% crash fault rate. In a bid to close this gap further, hosts might build a model of the distribution of its inputs. Hosts could then sample missing inputs from its posterior conditioned on successful messages. A natural first step would be to use a running mean as an estimator.
Appendix A Results on More Datasets, Comparison to FedVS
@@
In order to bolster the comparison to SplitNN, we evaluate the performance of DVFL on some benchmark vertically partitioned datasets. They are Credit Card Yeh & hui Lien (2009), Parkinson’s Sakar et al. (2019), CalTech-7 and Handwritten Markelle Kelly . Note that CalTech-7 is a subset of 7 classes CalTech 101 Li et al. (2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair. Both Handwritten and CalTech have 6 distinct “views”– features of distinct domains, e.g. spatial, frequency (Fourier), Gabor filter coefficients, etc.
@@
FedVS Li et al. (2023) is a state-of-the-art VFL algorithm which aims to tackle both privacy and fault tolerance. There is a round of secret shares so that each guest holds an encrypted version of a horizontal section of other guests’ data and model parameters. The shares are arranged in such a manner that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it. In order to maintain privacy despite this sharing of data, guests quantize their models and data and apply randomly mask. This means there is no longer enough information for a malicious guest to exactly reconstruct the data it receives from other guests. But, this obscurity is at the cost of model performance. Further, this algorithm is contingent on the guests using polynomial networks Livni et al. (2014a; b).
@@
In the following experiment, we compare our DVFL algorithm, SplitNN and FedVS. Our VFL and DVFL models are of the same depth as FedVS, but we use a smaller width. We also only use population normalization (as opposed to online batch and layer normalization). Therefore, the non-FedVS models have much fewer parameters and are simpler to train. We also take the liberty of using non-polynomial activations (Leaky ReLU). Model details and hyperparameters are in Appendix B. $|\mathcal{H}|=1$ for these experiments.
@@
Table 3: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions. Even on tabular datasets harder (especially for ANNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN. Despite using a smaller model, both outperform the results cited in the FedVS paper Li et al. (2023).
@@
FedVS Li et al. (2023) [^2] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
The feature splits and train-test splits for each dataset are identical to the FedVS paper Li et al. (2023). Further discussion about these results is available in Appendix C.
Appendix B Models, Hyperparameters and Implementation Details
B.1 Models
@@
B.1.1 MNIST
The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$, with Leaky ReLU activation (with negative half-plane slope = 0.01).
The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with a Leaky ReLU activation and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
@@ B.1.1 MNIST
Table 4: Ablation study of model width (particularly, the dimensions of the outputs which are communicated) on model performance with DVFL. Our experiments in the main paper use $W_{g}=320$ and $W_{h}=160$.
@@
B.1.2 VFL Datasets
Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in Appendix A. The model sizes for the guest $g_{i}$, host $h_{k}$ and owner $c$ are summarized in table 5, where $d_{i}$ is the input dimension of guest $g_{i}$.
@@ B.1.2 VFL Datasets
Credit Card $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ 10 Leaky ReLU Leaky ReLU
Parkinsons $\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor$ $\frac{\sum_{i}\lfloor{\frac{d_{i}}{4}}\rfloor}{4}$ Leaky ReLU ReLU
@@
Handwritten $\sum_{i}\lceil\frac{3d_{i}}{4}\rceil$ $3\frac{\sum_{i}\lceil\frac{3d_{i}}{4}\rceil}{4}$ Leaky ReLU ReLU
B.2 Fault Simulation
@@
    Communication Faults: During each communication round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$. We track the current status of each connection using flags.
@@
    In DVFL, we only update $\theta_{j}$ if the guest $g_{j}$ is alive. If it is a communication epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive. Similarly, in SplitNN, we only allow a guest to send a forward pass and receive gradients if it is alive.
@@ B.3 Other Hyperparameters
B.3.1 MNIST
For guests and DVFL hosts, we use an Adam optimizer with the default $\beta_{1}=0.9,\beta_{2}=0.999$. We use a learning rate of $10^{-3}$ and a weight decay of $10^{-5}$. For DVFL owners and SplitNN guests we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$. In general, guests were trained for 20 epochs, hosts for 40 and owners for 60. For SplitNN, the whole model was trained end-to-end with 60 epochs. For the experiments on limited intersecting datasets, owners/SplitNN hosts are trained for 160 epochs. In all cases, DVFL owner models and SplitNN host models were trained with early stopping.
B.3.2 VFL Datasets
Hyperparameters are presented in Table 6.
Table 6: Hyperparameters for the results presented in Appendix A.

### v2

@@ Appendix A DVFL for Split Neural Network Training
9:     $\theta_{h,i}\leftarrow\text{optim}_{h_{i}}(\theta_{h_{i}},\nabla_{\theta_{h_{i}}}\mathcal{L}_{h_{i}})$Update host model}
@@
Appendix B Results on More Datasets, Comparison to FedVS
@@
In order to bolster the comparison to SplitNN, we evaluate the performance of DVFL on some benchmark vertically partitioned datasets. They are Credit Card (Yeh & hui Lien, 2009), Parkinson’s (Sakar et al., 2019), CalTech-7 and Handwritten (Markelle Kelly, ). Note that CalTech-7 is a subset of 7 classes CalTech 101 (Li et al., 2022): Face, Motorbikes, Dolla-Bill, Garfield, Snoopy, Stop-Sign and Windsor-Chair. Both Handwritten and CalTech have 6 distinct “views”– features of distinct domains, e.g. spatial, frequency (Fourier), Gabor filter coefficients, etc.
@@
FedVS (Li et al., 2023) is a state-of-the-art VFL algorithm that aims to tackle both privacy and fault tolerance. There is a round of secret shares so that each guest holds an encrypted version of a horizontal section of other guests’ data and model parameters. The shares are arranged such that if $\leq L$ (a fixed system parameter) guests don’t send their outputs at the forward pass, there is enough redundant information in the other guests’ forward passes to account for it. In order to maintain privacy despite this sharing of data, guests quantize their models and data and apply randomly mask. This means there is no longer enough information for a malicious guest to exactly reconstruct the data it receives from other guests. But, this obscurity is at the cost of model performance. Further, this algorithm is contingent on the guests using polynomial networks (Livni et al., 2014a; b).
@@
In the following experiment, we compare our DVFL algorithm, SplitNN and FedVS. Our VFL and DVFL models are of the same depth as FedVS, but we use a smaller width. We also only use population normalization (as opposed to online batch and layer normalization). Therefore, the non-FedVS models have much fewer parameters and are simpler to train. We also take the liberty of using non-polynomial activations (Leaky ReLU). Model details and hyperparameters are in the Appendix. $|\mathcal{H}|=1$ for these experiments.
@@
Table 4: On a variety of typical vertically partitioned datasets, we see that DVFL performs comparably to SplitNN even under perfect conditions. Even on tabular datasets harder (especially for NNs) than MNIST such as Credit Card and Parkinson’s, DVFL model performance is comparable to SplitNN. Despite using a smaller model, both outperform the results cited in the FedVS paper (Li et al., 2023). DVFL achieves this while tolerating a broader class of faults and also providing complete immunity to gradient-based attacks.
@@
FedVS (Li et al., 2023) [^4] Results extracted from graphs. Raw figures were unavailable and the authors did not respond to our inquiries. 81.99 81.83 95.98 98.25
The feature splits and train-test splits for each dataset are identical to the FedVS paper (Li et al., 2023). Note that the datasets here are smaller than MNIST (30000 samples for credit card, 756 for Parkinsons, 1474 for MNIST, 2000 for handwritten). Observe that VFL and DVFL both outperform FedVS, which is quantized and constrained only to use polynomial networks (Livni et al., 2014a).
It is not meaningful to compare these methods under faults, since they are designed for different fault models. FedVS is completely tolerant to stragglers, particularly the case where stragglers are bounded in number. Like FedVS, this fault model does not affect DVFL, since guests and hosts train independently to each other. Conversely, our fault model might cause a catastrophic failure to FedVS: an arbitrary number of participants may crash, which is as if there is a number of stragglers greater than the predetermined system parameter.
Appendix C Other Failure Modes
Here, we have considered only crash faults. Other typical failure modes for distributed systems include arbitrary delays, incorrect order, and Byzantine faults.
Arbitrary Delays and Incorrect Order
As discussed in the previous section and 3, a line of work exists to partially tackle this problem. Typically, methods aimed at handling stragglers either demand a bound on the delay of stragglers or the number thereof. In return, methods such as FedVS (Li et al., 2023) can train with no performance loss. On the other hand, with bookkeeping like our VFL variants from the main paper, we may observe graceful degradation with stragglers and incorrect order (or no degradation at all in the case of wait, although that also requires a bounded delay).
Byzantine Faults
Byzantine faults are the most severe fault model, where messages may be corrupted. A possible future direction for DVFL is to incorporate tolerance to Byzantine faults. Observe that our SplitNN-buffer strategy is, in a sense, a Byzantine fault, and it is not performant. Byzantine faults are not particularly difficult for DVFL during the guest-host phase, since the learning target is not task-relevant. But some investigation is required for Byzantine faults during the owner learning phase, which cannot be outright detected as crash faults can.
Appendix D Models, Hyperparameters and Implementation Details
D.1 Models
@@
D.1.1 MNIST
Our batch sizes are 64.
The “guest” models have 2 layers, which squeeze the flattened $\frac{784}{|\mathcal{G}|}$-long input vector dimension to $\frac{400}{|\mathcal{G}|}$ via LeakyReLU, and then to $\frac{W_{g}}{|\mathcal{G}|}$ with ReLU activation.
The “host” models squeeze the $W_{g}$-dimensional input to $\frac{W_{g}+3W_{h}}{4}$ and then to $W_{h}$ via LeakyReLU, with Leaky ReLU activation (with negative half-plane slope = 0.01).
The “owner” model squeezes the $W_{h}\times|\mathcal{H}|$ dimensional input to $W_{h}$ and then to $40$ with Leaky ReLU activations and finally to a $10$-class prediction with an implicit softmax via the cross entropy loss function.
@@ D.1.1 MNIST
Table 5: Ablation study of model width (particularly, the dimensions of the outputs which are communicated) on model performance with DVFL. Our experiments in the main paper use $W_{g}=320$ and $W_{h}=160$. We run a hyperparameter sweep for each cell.
@@
D.1.2 VFL Datasets
Similar to MNIST, we use shallow multi-layer perceptrons for our results on vertically partitioned data in the Appendix. The model sizes for the guest $g_{i}$, host $h_{k}$ and owner $c$ are summarized in table 6, where $d_{i}$ is the input dimension of guest $g_{i}$.
@@ D.1.2 VFL Datasets
Credit Card $\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil$ 10 Leaky ReLU Leaky ReLU
Parkinsons $\sum_{i}\left\lfloor\frac{d_{i}}{4}\right\rfloor$ $\frac{\sum_{i}\left\lfloor\frac{d_{i}}{4}\right\rfloor}{4}$ Leaky ReLU ReLU
@@
Handwritten $\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil$ $3\frac{\sum_{i}\left\lceil\frac{3d_{i}}{4}\right\rceil}{4}$ Leaky ReLU ReLU
D.2 Fault Simulation
@@
    Connection Faults: During each connection round, for each pair $(g_{j},h_{i})$ we draw a sample from a uniform distribution with support $[0,1]$. We track the current status of each connection using flags.
@@
    In DVFL, we only update $\theta_{j}$ if the guest $g_{j}$ is alive. If it is a connection epoch, we only write $g_{j}$’s activations to $\mathbf{B}_{j,i}\forall i\in\{1,2,\dots,|\mathcal{H}|\}$ if the guest is alive. Similarly, in SplitNN, we only allow a guest to send a forward pass and receive gradients if it is alive.
@@ D.3 Other Hyperparameters
D.3.1 MNIST
For guests and DVFL hosts, we use an Adam optimizer with the default $\beta_{1}=0.9,\beta_{2}=0.999$. We use a learning rate of $2\times 10^{-3}$ (guest, host) and a weight decay of $10^{-5}$. For DVFL owners, we use SGD with a learning rate of $8\times 10^{-2}$. For SplitNN guests, we use stochastic gradient descent with a learning rate of $10^{-2}$ and momentum of $0.5$. In general, guests were trained for 20 epochs, hosts for 40 and owners for 60. For SplitNN, the whole model was trained end-to-end with 60 epochs. For the experiments on limited intersecting datasets, owners/SplitNN hosts are trained for 160 epochs. In all cases, DVFL owner models and SplitNN host models were trained with early stopping.
D.3.2 VFL Datasets
Hyperparameters are presented in Table LABEL:morehyper.
Table 7: Hyperparameters for the results presented in the Appendix.